OpenStudy (anonymous):

If the circumference of the circle below is 42 cm, what is the length of ?

OpenStudy (anonymous):

@Hero

hero (hero):

Something's missing bro.

OpenStudy (anonymous):

im looking for length of acr ab @Hero

OpenStudy (luigi0210):

@Danny_Boy do you know the formula?

hero (hero):

@Danny_Boy you need to make sure you're posting your problems completely and clearly.

hero (hero):

I think you mean you're looking for the length of arc ab.

OpenStudy (anonymous):

would it be length of arc AB \[\frac{ mAB }{ 360 } x \pi r ^{2}\]

hero (hero):

By the way: \[\frac{120}{360} = \frac{x}{42}\]

hero (hero):

Find x

OpenStudy (luigi0210):

@Hero you sure?

hero (hero):

@Luigi0210, do you think your formula is the only formula to solve it?

hero (hero):

You should try solving for x first

OpenStudy (anonymous):

what formula do i use :/

OpenStudy (luigi0210):

Just follow Hero for now

hero (hero):

Besides, @Luigi0210, your formula isn't right. Neither the area nor the radius is given.

OpenStudy (luigi0210):

I didn't give a formula Hero.

hero (hero):

Oops, my bad.

OpenStudy (anonymous):

should pi be in place of x :/

hero (hero):

I suppose neither of you are familiar with this proportion: \[\frac{\text{arc measure}}{360} = \frac{\text{arc length}}{\text{circumference}}\]

OpenStudy (anonymous):

im not :/

hero (hero):

Solve for x to get the arc length of AB

OpenStudy (anonymous):

ok so after dividing i get 0.33 :/

hero (hero):

How? You should have reduced 120/360 to get 1/3 Then you should have \[\frac{1}{3} = \frac{x}{42}\] From there, you cross multiply and solve for x.

OpenStudy (anonymous):

ohhhh i c 42=3x :/

OpenStudy (anonymous):

after that i divide right...... and i get 0.07 :/

hero (hero):

You might need to work on your division skills.

hero (hero):

\[\frac{42}{3} = 14\]

OpenStudy (anonymous):

ahhh true i did some thing wrong :P

hero (hero):

Calc fail

OpenStudy (anonymous):

lol ahhh i got it now thank u :D

hero (hero):

Bro, you should fix that problem you have before you mess up on a test.

OpenStudy (anonymous):

yeah i will have tio go over and fix every thing thanks though